A sample of gas has a pressure of 2.9 atm and a temperature of 435K.  What will be the new temperature if the pressure is

increased to 4.0 atm?


1
Expert's answer
2021-04-05T07:40:27-0400

Using Gay-lussac law,

P1/T1 = P2/T2

2.9/435 = 4.0/T2

T2 = 600 K

So, New temperature is = [600] K